Dale is 4 years older than Sue. 5 years ago the sum of their ages was 64. How old are they now?
Answer by jorel1380(3719)

Let X be Sue's age. Dale is 4 years older than Sue, or X+4. If the sum of their ages 5 years ago was 64, then:
(X-5)+((X+4)-5)=64
2X-6=64
2X=70
X=35
Therefore, Sue is now 35, Dale is 39.
35+39-10=64.